OpenStudy (anonymous):

x + 2 3-x 2 ----- + ---- = --- 15 5 3 Solve.

OpenStudy (anonymous):

take LCM of the denominators on left hand side and then solve it....hope u understand it

OpenStudy (anonymous):

15 will be the LCM

OpenStudy (anonymous):

x+2+3(3-x)/15=2/3 x+2+9-3x/15=2/3 -2x+11/15=2/3 -2x+11=2/3*15 -2x+11=10 -2x=10-11 -2x=-1 x=1/2

OpenStudy (anonymous):

cool, thanks

OpenStudy (anonymous):

@unknown&confused hope you understand it

OpenStudy (anonymous):

u r welcum =)

hero (hero):

You don't need to use LCM to solve this. Just multiply the second fraction by 3/3

hero (hero):

Then add the fractions on the left side, then cross multiply